# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 2.2.3 Conversational Engagements in Education
Early LLM-based tutors often fell into a simple question-answer pattern, providing direct answers and explanations that made students passive. To address this, researchers have developed Socratic and guided approaches that engage learners in multi-turn dialogue. SocraticLM [136] is a notable example, proposing a “thought-provoking” teaching paradigm in which the tutor asks open-ended questions and prompts the student’s reasoning instead of giving away solutions. SocraticLM was trained on SocraTeach, a new dataset of 35k multi-turn dialogues where a simulated teacher guides students with diverse cognitive states through math problems. Similarly, Kargupta et al. (2024) [68] tackle the “answer-too-direct” issue in the coding domain with TreeInstruct, an LLM-based tutor that plans a hierarchy of questions to help students debug code. TreeInstruct models the student’s knowledge state and asks targeted questions for each error, effectively guiding learners to independently correct mistakes. It achieved state-of-the-art results on code debugging benchmarks and demonstrated in a user study that students could fix bugs with minimal direct hints.
Beyond specific algorithms, others have looked at controlling LLM tutors via high-level pedagogical strategies. For example, StratL [137] introduces a pedagogical steering framework: instead of letting an LLM tutor freely generate replies, StratL optimizes prompts to make the tutor follow a predefined teaching plan represented as a graph. In a case study on Productive Failure (an educational strategy where students first struggle with problems before instruction), StratL successfully steered an LLM tutor to withhold answers and encourage productive trial-and-error. Notably, in a field experiment with 17 high school students, the LLM tutor guided by StratL adhered to the desired strategy and helped students discover solutions on their own. These efforts collectively show a shift from straightforward Q&A to multi-turn, strategy-aware dialogue, making LLM tutors behave more like human teachers who ask, hint, and adapt rather than just tell.
@@
Traditional intelligent tutoring systems often struggled with a “one size fits all” design that failed to accommodate individual learners. In contrast, many new approaches use conversational turns to tailor teaching to student needs. PACE (PersonAlized Conversational tutoring agEnt) [138], introduced by Liu et al. (2025) under the motto “One Size Doesn’t Fit All,” explicitly models student learning styles and personas to personalize the dialogue. PACE’s multi-turn tutor adapts its explanations and questions according to the Felder–Silverman learning style model – for instance, giving concrete examples for sensory learners versus abstract prompts for intuitive learners. It also incorporates the Socratic method (asking guiding questions instead of lecturing) to stimulate critical thinking. To enable this, a new dataset of personalized tutoring dialogues was created, simulating students with diverse backgrounds and personalities and pairing them with an LLM tutor.
Moreover, LLM tutors can track a student’s step-by-step reasoning in multi-turn exchanges and provide targeted help at the right moment. This ability to model the student within the conversation marks a key advantage of multi-turn tutoring. An illustrative real-world deployment is JeepyTA [139], a GPT-based virtual teaching assistant used in an online course forum. JeepyTA monitors student questions on course material and responds in seconds with context-specific help, adapting its style to the informal, conversational tone of forum discussions. It distinguishes logistical queries from conceptual ones and provides hints or explanations accordingly. By continuously engaging with students’ follow-up questions, such an always-on conversational TA exemplifies how personalization and instant adaptivity can scale via LLMs.
@@
Recent research further explores ways to deepen assessments of LLM tutoring capabilities, particularly regarding their subject expertise and pedagogical effectiveness in mathematics. An influential early example is the work of Macina et al. (2023) [53], who introduced MathDial to systematically evaluate LLMs’ abilities as tutors, emphasizing faithful and equitable teaching. Their approach involved collecting high-quality, teacher–student dialogues through human–LLM interactions, with InstructGPT simulating student behaviors—including common misconceptions—and expert annotators assuming the teacher role. Diverse student responses were elicited using temperature sampling, prompting the LLM to generate realistic errors. Human teachers then employed scaffolding strategies to guide the simulated students toward solutions. Dialogue quality was ensured through rigorous human annotation. Evaluation metrics included the simulated student’s success rate and the telling@k score, indicating how frequently teachers prematurely revealed answers. Models fine-tuned on the MathDial dataset demonstrated improved student outcomes by emphasizing hints and prompting over direct answers. Similarly, Ding et al. (2024) [140] introduced SocraticLLM, a knowledge-enhanced model emulating Socratic questioning to foster critical thinking and self-discovery. To support this method, they developed the publicly available SocraticMATH dataset, containing structured Socratic dialogues that cover 513 primary-school math topics. Another benchmark, MathTutorBench by Macina et al. (2025) [141], specifically targets one-on-one math tutor–student dialogues, assessing LLM tutors across multiple dimensions, including Math Expertise (accuracy in problem-solving), Student Understanding (ability to diagnose and correct misconceptions), and Teacher Response Quality (effectiveness in providing hints and Socratic guidance).
@@
Most recently, retrieval-augmented generation (RAG) techniques have further advanced the field. For instance, Feng et al. (2024) [142] developed CourseAssist, a system that grounds tutor-generated responses explicitly in course syllabi and lecture notes, ensuring alignment with instructor expectations. Levonian et al. (2025) [143] explored alignment methods to enhance generative tutors’ outputs, making responses safer, pedagogically appropriate, and closely relevant to student queries, significantly improving multi-turn algebra tutoring dialogues. Expanding upon this theme, Scarlatos et al. (2025) [144] argue for the necessity of adaptive AI tutors that subtly adjust responses across dialogues, strategically guiding students toward correct understanding without explicit intervention. They operationalized this idea by generating multiple potential tutor replies at each conversational turn and then assessing them with a student model—which predicts students’ subsequent correctness—and a pedagogical evaluator. By ranking tutor responses based on these dual criteria, they successfully fine-tuned a new tutor model through direct preference optimization, thereby rewarding interactions that consistently promoted student learning.
@@
The emerging trend in conversational tutors involves integrating research insights directly into classroom practice, facilitated by the increasing availability of large language models tailored specifically for education. Major industry players, including Google with LearnLM [145] and Anthropic with Claude for Education [146], have actively embraced this approach, developing specialized educational versions of their flagship AI models. These initiatives reflect a convergence between academic research and industry applications, emphasizing personalized, adaptive tutoring methods grounded in evidence-based dialogue strategies and instructional principles.
@@
Large language models are increasingly leveraged to generate automated feedback and grading, aiming to support instructors with large-scale assessments [147]. Similarly, in open-ended writing tasks, LLMs can produce fluent and plausible comments that appear insightful, yet often include content not grounded in the student’s work [148]. This lack of faithfulness (e.g. fabricated critiques or irrelevant suggestions) is a critical limitation, as unfaithful feedback can mislead or confuse learners. Improving the accuracy and alignment of feedback with students’ actual mistakes has therefore become a central research focus.
@@
To address these issues, recent work has introduced strategies to make LLM feedback more interactive, adaptive, and pedagogically grounded. One approach is to incorporate a verification step before feedback delivery: Daheim et al. (2024) have an LLM “verifier” analyze the student’s reasoning step-by-step to pinpoint errors, which then guides a tutor model to give targeted hints [149]. This stepwise verification significantly reduces hallucinated advice and yields feedback more precisely tailored to the student’s misunderstanding. Another line of research optimizes feedback through learning from human pedagogical preferences. Scarlatos et al. (2024) [150] propose a rubric-based evaluation of feedback quality (checking for correctness, encouragement, misconception-addressing, etc.), use GPT-4 to label LLM outputs along these criteria, and then apply reinforcement learning to tune the model . The result is an LLM that produces feedback with measurably higher factual correctness and better alignment to effective tutoring practices. There are even efforts to “close the loop” by having the LLM anticipate student revisions: Nair et al. (2024) [151] introduce a system where the model generates feedback, simulates how a student would revise their essay in response, and iteratively refines its feedback to maximize the improvement between drafts . This led to greater actual improvements in student writing compared to static feedback, and the optimized feedback exhibited enhanced pedagogical qualities.
@@
Studies are also examining how well LLM feedback aligns with real classroom needs and how students/instructors perceive it. Initial deployments of LLM-based feedback systems in university courses show mixed but encouraging results. In a graduate-level computer science class, an automated feedback tool generated paragraph-level comments on project reports, which students found generally helpful for improving their work [152]. The course instructor noted, however, that the AI’s comments did not always match the assignment’s pedagogical objectives or emphasis [153]. In fact, the instructor preferred to use the LLM’s output as a draft – a starting point that the instructor would edit – rather than sending it raw to students. This highlights an important limitation: current LLMs may need human oversight to ensure feedback is consistent with the teacher’s goals and curriculum. On the other hand, domain-focused applications of LLM feedback have shown clear benefits when carefully integrated. Riazi and Rooshenas developed an LLM-driven tutor for a databases course that could analyze a student’s entity-relationship diagram and generate detailed, context-specific critiques and follow-up questions [154]. Likewise, in medical education, LLMs have been used to generate explanatory feedback for answers to medical board-style multiple-choice questions. Experts evaluating such feedback found it relevant and useful – not a replacement for human feedback, but a valuable supplement to the usual numeric scores students receive [152].
@@
Efforts like Lohr et al. (2025) [155]’s, which prompt LLMs to produce specific types of feedback from established educational taxonomies (e.g. an error-identification vs. a hint vs. an elaboration), further illustrate the push toward more controlled and purposefully designed feedback messages. In addition to feedback itself, researchers are beginning to evaluate LLMs on related teaching skills such as asking good questions – for instance, the Dr.Academy benchmark assesses whether LLMs can generate high-quality, higher-order questions in line with Bloom’s taxonomy, finding that models like GPT-4 already show strong capability in formulating deep conceptual questions. Together, these advances show a clear developmental trajectory: from basic feedback generation that often wandered off-target, to increasingly faithful, adaptive, and pedagogically-aware feedback loops facilitated by LLMs’ multi-turn interaction capacity.
@@
In parallel with feedback generation, researchers have started leveraging LLMs to grade student work and provide evaluative judgments (scores, ratings, or rubric-based assessments). Automatic grading by AI is not entirely new, but LLMs offer a unified, flexible approach that can handle open-ended responses more like a human grader. Recent studies suggest that, for certain types of assignments, LLM graders can approach human-level performance in both consistency and accuracy. Capdehourat et al. (2025) [156] explored LLMs for scoring short free-response questions in Spanish, a scenario involving language complexity beyond the typical English-centric datasets. They found that state-of-the-art models (including GPT-4 and advanced open-source LLMs) could predict expert graders’ scores with over 95% accuracy in a three-tier grading scale, and even 98% accuracy on simpler right/wrong judgments.
@@
Another study compared ChatGPT directly against university instructors for grading full exam papers in higher education. Out of 463 Master’s-level exam responses graded, about 70% of the AI’s assigned scores fell within a 10% margin of the human-given score, and  31% were within a 5% margin [157]. Teachers involved in the experiment expressed surprise at how closely ChatGPT’s evaluations matched their own in these exams. However, important discrepancies were observed. The AI grader tended to be more conservative, avoiding very high or very low scores on individual questions.
@@
In the domain of essay scoring, which demands understanding of content, organization, style, and often providing feedback, LLMs still struggle to meet human-level nuance. Kostić et al. (2024) [158] conducted a case study using GPT-4 to evaluate German-language student essays at a business school . The LLM could generate a score and some comments, but it often failed to apply the rubric criteria consistently and lacked the depth of feedback that human lecturers provided. Complex aspects of writing quality (critical analysis, creativity, etc.) proved difficult for the model to judge correctly, highlighting a gap between what current LLMs can do and the “nuanced requirements” of real essay evaluation.
@@
Scenario Simulation Recent advances have explored scenario simulation as a means to leverage LLMs for enacting multi-turn educational interactions between virtual teachers and students. Early work focused on using LLM-simulated student profiles to evaluate learning materials. For example, Generative Students introduced a prompt-based architecture (grounded in the Knowledge-Learning-Instruction framework) to instantiate diverse student profiles defined by mastered vs. confused knowledge components [159]. Each simulated student (powered by GPT-4) answered multiple-choice questions, producing responses that aligned with its knowledge profile. Notably, these generative students exhibited answer patterns highly correlated with real student performance, correctly flagging many of the same difficult questions. This result suggested that realistic virtual learners can serve as proxies for human students in content evaluation, allowing instructors to identify problematic questions before deployment.
@@
Subsequent research broadened the fidelity of simulated students by incorporating individual differences in ability and personality. Liu et al. (2024) [160] developed a personality-aware simulation framework that enriches student profiles with both cognitive level (e.g. language proficiency) and noncognitive traits (e.g. conscientiousness). In a language tutoring scenario, an LLM could then produce diverse student utterances consistent with a given persona, which in turn successfully triggered the tutor’s adaptive scaffolding strategies. Building on this idea, Jin et al. (2024) [161] introduced TeachTune, a system enabling teachers to test their pedagogical conversational agents (PCAs) against diverse simulated students. Teachers specify a student’s presumed prior knowledge and motivation, and an LLM-driven student agent engages in a multi-turn chat with the PCA. This automated student–teacher dialogue reveals how well the PCA adapts its explanations and feedback to different learner needs, going beyond single-turn Q&A tests. The TeachTune pipeline ensured that each simulated student’s behavior remained faithful to its profile, with measured deviations under 5–10%. These efforts highlight that richly modeled virtual students can support teacher agents and tutoring systems by surfacing potential shortcomings in adaptive instruction in a cost-effective manner.
@@
Researchers have also scaled scenario simulation to multi-party settings. Zhang et al. (2024) [162] proposed SimClass, a framework in which multiple LLM-based agents assume typical classroom roles. A novel class-level control mechanism orchestrates the agents’ turn-taking and topic flow to emulate a live classroom lesson. In user trials with real students, SimClass was able to simulate dynamic classroom interactions featuring both teacher–student exchanges and student–student discussions. The emergent group behavior was strikingly human-like – the student agents would ask and answer each other’s questions and collaboratively debate topics, creating an enlivened atmosphere. These collective simulations improved the human participant’s learning experience by maintaining engagement and peer-like dialogue support. The success of SimClass demonstrates that LLMs can collectively model complex social dynamics of a classroom, opening the door to virtual class rehearsals and large-scale peer interaction scenarios.
@@
Later on, scenario simulation has been extended to address specialized pedagogical needs for both learners and instructors. To better support students with low academic performance or poor metacognitive skills, Li et al. (2025) [163] devised a pipeline for generating struggling student agents and evaluating their realism. This approach automatically creates a spectrum of student profiles with varying learning deficiencies and filters them through a two-round LLM evaluation (validated by human experts) to ensure the simulated learning struggles are authentic. By assembling a set of high-fidelity “at-risk” student agents, educators and intelligent tutoring systems can safely experiment with interventions to foster self-regulation and reflective thinking, without ethical concerns of testing on real students. On the instructor side, Hu et al. (2025) [164] explored using LLM-based teaching simulation to improve lesson planning. In their method, an LLM is prompted to play out a full classroom lesson based on a teacher’s draft plan – simulating the teacher’s instruction and the students’ reactions – and then to generate a reflective critique of that session. The insights from this simulated teacher–student interaction are used to iteratively refine the lesson plan. Additionally, Wang et al. (2024) [165] proposed Book2Dial, a framework to automatically generate synthetic teacher-student dialogues from textbook content to address data scarcity in developing educational chatbots. Three dialogue-generation approaches—Multi-turn QG-QA, Dialogue Inpainting, and LLM-based Role-Playing—are introduced and evaluated using automated metrics (e.g., coherence, answerability, factual consistency) and human judgment (specificity). The results demonstrate that LLM-based Role-Playing performs best, highlighting a cost-effective method for chatbot training in educational domains.

### v5

@@ 2.2.3 Conversational Engagements in Education
Early LLM-based tutors often fell into a simple question-answer pattern, providing direct answers and explanations that made students passive. To address this, researchers have developed Socratic and guided approaches that engage learners in multi-turn dialogue. SocraticLM (Liu et al., 2024a) is a notable example, proposing a “thought-provoking” teaching paradigm in which the tutor asks open-ended questions and prompts the student’s reasoning instead of giving away solutions. SocraticLM was trained on SocraTeach, a new dataset of 35k multi-turn dialogues where a simulated teacher guides students with diverse cognitive states through math problems. Similarly, Kargupta et al. (2024) tackle the “answer-too-direct” issue in the coding domain with TreeInstruct, an LLM-based tutor that plans a hierarchy of questions to help students debug code. TreeInstruct models the student’s knowledge state and asks targeted questions for each error, effectively guiding learners to independently correct mistakes. It achieved state-of-the-art results on code debugging benchmarks and demonstrated in a user study that students could fix bugs with minimal direct hints.
Beyond specific algorithms, others have explored controlling LLM tutors through high-level pedagogical strategies. StratL (Puech et al., 2025) introduces a pedagogical steering framework that optimizes prompts to guide an LLM tutor along a predefined teaching plan represented as a graph, rather than allowing unconstrained generation. Applied to Productive Failure—a strategy where students wrestle with problems before receiving instruction—StratL successfully steers the tutor to withhold answers and encourage trial-and-error. In a field experiment with 17 high school students, the StratL-guided tutor adhered to the target strategy and facilitated student-driven discovery. Collectively, these efforts mark a shift from straightforward Q&A toward multi-turn, strategy-aware dialogue, bringing LLM tutors closer to human teachers who ask, hint, and adapt rather than simply tell.
@@
Traditional intelligent tutoring systems often suffered from a "one-size-fits-all" design that poorly accommodated individual learners. PACE (PersonAlized Conversational tutoring agEnt) (Liu et al., 2025a) directly addresses this limitation by explicitly modeling student learning styles and personas. Built on the Felder–Silverman learning style model, PACE’s multi-turn tutor adapts its explanations accordingly—providing concrete examples for sensory learners and abstract prompts for intuitive ones—while employing Socratic questioning to stimulate critical thinking. To support this, the authors construct a new dataset of personalized tutoring dialogues simulating students with diverse backgrounds and personalities interacting with an LLM tutor.
Moreover, LLM tutors can track a student’s step-by-step reasoning in multi-turn exchanges and provide targeted help at the right moment. This ability to model the student within the conversation marks a key advantage of multi-turn tutoring. An illustrative real-world deployment is JeepyTA (Liu et al., 2025d), a GPT-based virtual teaching assistant used in an online course forum. JeepyTA monitors student questions on course material and responds in seconds with context-specific help, adapting its style to the informal, conversational tone of forum discussions. It distinguishes logistical queries from conceptual ones and provides hints or explanations accordingly. By continuously engaging with students’ follow-up questions, such an always-on conversational TA exemplifies how personalization and instant adaptivity can scale via LLMs.
@@
Recent research further explores ways to deepen assessments of LLM tutoring capabilities, particularly regarding their subject expertise and pedagogical effectiveness in mathematics. An influential early example is the work of Macina et al. (2023), who introduced MathDial to systematically evaluate LLMs’ abilities as tutors, emphasizing faithful and equitable teaching. Their approach involved collecting high-quality, teacher–student dialogues through human–LLM interactions, with InstructGPT simulating student behaviors—including common misconceptions—and expert annotators assuming the teacher role. Diverse student responses were elicited using temperature sampling, prompting the LLM to generate realistic errors. Human teachers then employed scaffolding strategies to guide the simulated students toward solutions. Dialogue quality was ensured through rigorous human annotation. Evaluation metrics included the simulated student’s success rate and the telling@k score, indicating how frequently teachers prematurely revealed answers. Models fine-tuned on the MathDial dataset demonstrated improved student outcomes by emphasizing hints and prompting over direct answers. Similarly, Ding et al. (2024) introduce SocraticLLM, a knowledge-enhanced model emulating Socratic questioning to foster critical thinking and self-discovery. To support this method, they developed the publicly available SocraticMATH dataset, containing structured Socratic dialogues that cover 513 primary-school math topics. Another benchmark, MathTutorBench by Macina et al. (2025), specifically targets one-on-one math tutor–student dialogues, assessing LLM tutors across multiple dimensions, including Math Expertise (accuracy in problem-solving), Student Understanding (ability to diagnose and correct misconceptions), and Teacher Response Quality (effectiveness in providing hints and Socratic guidance).
@@
Most recently, retrieval-augmented generation (RAG) techniques have further advanced the field. For instance, Feng et al. (2024) develop CourseAssist, a system that grounds tutor-generated responses explicitly in course syllabi and lecture notes, ensuring alignment with instructor expectations. Levonian et al. (2025) explore alignment methods to enhance generative tutors’ outputs, making responses safer, pedagogically appropriate, and closely relevant to student queries, significantly improving multi-turn algebra tutoring dialogues. Expanding upon this theme, Scarlatos et al. (2025) argue for the necessity of adaptive AI tutors that subtly adjust responses across dialogues, strategically guiding students toward correct understanding without explicit intervention. They operationalize this idea by generating multiple potential tutor replies at each conversational turn and then assessing them with a student model—which predicts students’ subsequent correctness—and a pedagogical evaluator. By ranking tutor responses based on these dual criteria, they successfully fine-tune a new tutor model through direct preference optimization, thereby rewarding interactions that consistently promoted student learning.
@@
The emerging trend in conversational tutors involves integrating research insights directly into classroom practice, facilitated by the increasing availability of large language models tailored specifically for education. Major industry players, including Google with LearnLM (Wiggers, 2024) and Anthropic with Claude for Education (Anthropic, 2025), have actively embraced this approach, developing specialized educational versions of their flagship AI models. These initiatives reflect a convergence between academic research and industry applications, emphasizing personalized, adaptive tutoring methods grounded in evidence-based dialogue strategies and instructional principles.
@@
Large language models are increasingly leveraged to generate automated feedback and grading, aiming to support instructors with large-scale assessments (Silva and Costa, 2025). Similarly, in open-ended writing tasks, LLMs can produce fluent and plausible comments that appear insightful, yet often include content not grounded in the student’s work (Jia et al., 2024b). This lack of faithfulness (e.g. fabricated critiques or irrelevant suggestions) is a critical limitation, as unfaithful feedback can mislead or confuse learners. Improving the accuracy and alignment of feedback with students’ actual mistakes has therefore become a central research focus.
@@
To address these limitations, recent work has introduced strategies to make LLM feedback more interactive, adaptive, and pedagogically grounded. Daheim et al. (2024) incorporate a verification step in which an LLM first analyzes student reasoning step-by-step to pinpoint errors, then guides a tutor model to deliver targeted hints—significantly reducing hallucinated advice and improving alignment to the student’s actual misunderstanding. Scarlatos et al. (2024) instead optimize feedback quality through human pedagogical preferences: they define a rubric covering correctness, encouragement, and misconception-addressing, use GPT-4 to label outputs along these criteria, and apply reinforcement learning to fine-tune the model, yielding measurably higher factual correctness and better tutoring alignment. Going further, Nair et al. (2024) close the feedback loop by having the model simulate student revisions in response to its own feedback and iteratively refine that feedback to maximize draft improvement—producing greater actual writing gains and enhanced pedagogical qualities compared to static feedback.
@@
Studies are also examining how well LLM feedback aligns with real classroom needs and how students and instructors perceive it. Initial deployments show mixed but encouraging results. In a graduate computer science course, an automated tool generated paragraph-level comments on project reports that students found broadly helpful (Tomova et al., 2024), yet the instructor observed that the AI’s comments did not always reflect the assignment’s pedagogical objectives—preferring to treat LLM output as a draft to edit rather than send directly to students (Jia et al., 2024a). This underscores a recurring limitation: LLM feedback often requires human oversight to remain consistent with instructional goals. Domain-focused deployments, however, demonstrate clearer benefits when carefully integrated. Riazi and Rooshenas (2025) develop an LLM-driven tutor for a databases course capable of analyzing student entity-relationship diagrams and generating detailed, context-specific critiques and follow-up questions. Similarly, in medical education, LLMs have been used to produce explanatory feedback for board-style multiple-choice questions; expert evaluators found such feedback relevant and useful—not a replacement for human judgment, but a meaningful supplement to the numeric scores students typically receive (Tomova et al., 2024).
@@
Efforts like Lohr et al. (2025)’s, which prompt LLMs to produce specific types of feedback from established educational taxonomies (e.g. an error-identification vs. a hint vs. an elaboration), further illustrate the push toward more controlled and purposefully designed feedback messages. In addition to feedback itself, researchers are beginning to evaluate LLMs on related teaching skills such as asking good questions – for instance, the Dr.Academy benchmark assesses whether LLMs can generate high-quality, higher-order questions in line with Bloom’s taxonomy, finding that models like GPT-4 already show strong capability in formulating deep conceptual questions. Together, these advances show a clear developmental trajectory: from basic feedback generation that often wandered off-target, to increasingly faithful, adaptive, and pedagogically-aware feedback loops facilitated by LLMs’ multi-turn interaction capacity.
@@
In parallel with feedback generation, researchers have started leveraging LLMs to grade student work and provide evaluative judgments (scores, ratings, or rubric-based assessments). Automatic grading by AI is not entirely new, but LLMs offer a unified, flexible approach that can handle open-ended responses more like a human grader. Recent studies suggest that, for certain types of assignments, LLM graders can approach human-level performance in both consistency and accuracy. Capdehourat et al. (2025) explore LLMs for scoring short free-response questions in Spanish, a scenario involving language complexity beyond the typical English-centric datasets. They found that state-of-the-art models (including GPT-4 and advanced open-source LLMs) could predict expert graders’ scores with over 95% accuracy in a three-tier grading scale, and even 98% accuracy on simpler right/wrong judgments.
@@
Another study compared ChatGPT directly against university instructors for grading full exam papers in higher education. Out of 463 Master’s-level exam responses graded, about 70% of the AI’s assigned scores fell within a 10% margin of the human-given score, and  31% were within a 5% margin (Flodén, 2025). Teachers involved in the experiment expressed surprise at how closely ChatGPT’s evaluations matched their own in these exams. However, important discrepancies were observed. The AI grader tended to be more conservative, avoiding very high or very low scores on individual questions.
@@
In the domain of essay scoring, which demands understanding of content, organization, style, and often providing feedback, LLMs still struggle to meet human-level nuance. Kostic et al. (2024) conduct a case study using GPT-4 to evaluate German-language student essays at a business school . The LLM could generate a score and some comments, but it often failed to apply the rubric criteria consistently and lacked the depth of feedback that human lecturers provided. Complex aspects of writing quality (critical analysis, creativity, etc.) proved difficult for the model to judge correctly, highlighting a gap between what current LLMs can do and the “nuanced requirements” of real essay evaluation.
@@
Scenario Simulation Recent advances have explored scenario simulation as a means to leverage LLMs for enacting multi-turn educational interactions between virtual teachers and students. Early work focused on using LLM-simulated student profiles to evaluate learning materials. For example, Generative Students introduced a prompt-based architecture (grounded in the Knowledge-Learning-Instruction framework) to instantiate diverse student profiles defined by mastered vs. confused knowledge components (Lu and Wang, 2024). Each simulated student (powered by GPT-4) answered multiple-choice questions, producing responses that aligned with its knowledge profile. Notably, these generative students exhibited answer patterns highly correlated with real student performance, correctly flagging many of the same difficult questions. This result suggested that realistic virtual learners can serve as proxies for human students in content evaluation, allowing instructors to identify problematic questions before deployment.
@@
Subsequent research broadened the fidelity of simulated students by incorporating individual differences in ability and personality. Liu et al. (2024f) develop a personality-aware simulation framework that enriches student profiles with both cognitive level (e.g. language proficiency) and noncognitive traits (e.g. conscientiousness). In a language tutoring scenario, an LLM could then produce diverse student utterances consistent with a given persona, which in turn successfully triggered the tutor’s adaptive scaffolding strategies. Building on this idea, Jin et al. (2025) introduce TeachTune, a system enabling teachers to test their pedagogical conversational agents (PCAs) against diverse simulated students. Teachers specify a student’s presumed prior knowledge and motivation, and an LLM-driven student agent engages in a multi-turn chat with the PCA. This automated student–teacher dialogue reveals how well the PCA adapts its explanations and feedback to different learner needs, going beyond single-turn Q&A tests. The TeachTune pipeline ensured that each simulated student’s behavior remained faithful to its profile, with measured deviations under 5–10%. These efforts highlight that richly modeled virtual students can support teacher agents and tutoring systems by surfacing potential shortcomings in adaptive instruction in a cost-effective manner.
@@
Researchers have also scaled scenario simulation to multi-party settings. Zhang et al. (2025b) propose SimClass, a framework in which multiple LLM-based agents assume typical classroom roles. A novel class-level control mechanism orchestrates the agents’ turn-taking and topic flow to emulate a live classroom lesson. In user trials with real students, SimClass was able to simulate dynamic classroom interactions featuring both teacher–student exchanges and student–student discussions. The emergent group behavior was strikingly human-like – the student agents would ask and answer each other’s questions and collaboratively debate topics, creating an enlivened atmosphere. These collective simulations improved the human participant’s learning experience by maintaining engagement and peer-like dialogue support. The success of SimClass demonstrates that LLMs can collectively model complex social dynamics of a classroom, opening the door to virtual class rehearsals and large-scale peer interaction scenarios.
@@
Scenario simulation has since been extended to specialized pedagogical needs for both learners and instructors. To better support low-performing or metacognitively weak students, Li et al. (2025a) devise a pipeline that automatically generates struggling student agents with diverse learning deficiencies, then filters them through a two-round LLM evaluation—validated by human experts—to ensure authenticity. The resulting high-fidelity "at-risk" agents allow educators and tutoring systems to safely experiment with interventions targeting self-regulation and reflective thinking, without ethical concerns of testing on real students. On the instructor side, Hu et al. (2025a) use LLM-based simulation to improve lesson planning: an LLM plays out a full classroom lesson from a teacher’s draft plan, simulating both instruction and student reactions, then generates a reflective critique. These simulated insights are used to iteratively refine the lesson plan before deployment. Additionally, Wang et al. (2024d) propose Book2Dial, a framework to automatically generate synthetic teacher-student dialogues from textbook content to address data scarcity in developing educational chatbots. Three dialogue-generation approaches—Multi-turn QG-QA, Dialogue Inpainting, and LLM-based Role-Playing—are introduced and evaluated using automated metrics (e.g., coherence, answerability, factual consistency) and human judgment (specificity). The results demonstrate that LLM-based Role-Playing performs best, highlighting a cost-effective method for chatbot training in educational domains.